OpenStudy (anonymous):

write the equation -4x=2y=1 in slope-intercept form. give the slope and the y-intercept.

OpenStudy (p0sitr0n):

how is -4x = 2y =1

OpenStudy (anonymous):

sorry -4x+2y=1

OpenStudy (p0sitr0n):

2y = 1+4x y = 2x +0.5 2 = slope y intercept = x =0 y = 0.5